# Nested Sampling for ARIMA Model Selection in Astronomical Time-Series Analysis

### v1

@@ Acknowledgements
The sunspots time series data used in this work were obtained from the sunspots dataset provided in the statsmodels library (credit: SIDC, RWC Belgium, World Data Center for the Sunspot Index, Royal Observatory of Belgium, 1700-2008). The Kepler and TESS photometric lightcurves were obtained and processed using the lightkurve (Lightkurve Collaboration et al., 2018) Python package. All the data used in this analysis, including the relevant nested sampling chains, is available at (Naik, 2025). We also include a python notebook to generate the plots and figures presented in this paper.
@@ Appendix A Comparison of Forecasting Performance
The forecasting performance of ARIMA(9,1)(9,1) and ARIMA(3,3)(3,3) are compared in Figure 18.
Figure 18: Comparison of posterior predictive forecasts for two ARIMA models applied to the yearly sunspots-number time series. Panels (a) and (b) show the posterior predictive distributions of the ARIMA(9,1)(9,1) and ARIMA(3,3)(3,3) models, respectively with shaded contours denoting the $1\sigma$$2\sigma$ and $3\sigma$ credible regions of the predictive posterior distribution $P(\hat{y}_{t}|t,D_{t})$.
ARIMA(9,1)(9,1) achieves a markedly better overall forecasting performance as compared to ARIMA(3,3)(3,3). It shows the expected growth of predictive uncertainty with time, which is characteristic of long-term ARIMA forecasting. ARIMA(9,1)(9,1) forecasts also yielded much lower values for the Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) (see Table 6), thereby confirming and validating the results of our model selection methodology.

### v2

@@ Acknowledgements
The sunspots time series data used in this work were obtained from the sunspots dataset provided in the statsmodels library (credit: SIDC, RWC Belgium, World Data Center for the Sunspot Index, Royal Observatory of Belgium, 1700-2008). The Kepler and TESS photometric light curves were obtained and processed using the lightkurve (Lightkurve Collaboration et al., 2018) Python package. All the data used in this analysis, including the relevant nested sampling chains, is available at (Naik, 2026). We also include a python notebook to generate the plots and figures presented in this paper.
@@ Appendix A Comparison of Forecasting Performance
The forecasting performance of ARIMA(9,0,1)(9,0,1) and ARIMA(3,0,3)(3,0,3) are compared in Figure 31.
Figure 31: Comparison of posterior predictive forecasts for two ARIMA models applied to the yearly sunspot number time series. Panels (a) and (b) show the posterior predictive distributions of the ARIMA(9,0,1)(9,0,1) and ARIMA(3,0,3)(3,0,3) models, respectively with shaded contours denoting the $1\sigma$$2\sigma$ and $3\sigma$ credible regions of the predictive posterior distribution $P(\hat{y}_{t}|t,D_{t})$.
ARIMA(9,0,1)(9,0,1) achieves a markedly better overall forecasting performance compared to ARIMA(3,0,3)(3,0,3). It shows the expected growth of predictive uncertainty with time, which is characteristic of long-term ARIMA forecasting. ARIMA(9,0,1)(9,0,1) forecasts also yielded much lower values for the Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) (see Table 13), thereby confirming and validating the results of our model selection methodology.
@@
Appendix B Computational Performance
Table 14 reports wall-clock runtimes for ARIMA grid searches of increasing size, evaluated on a synthetic time series of $N=1000$ data points. We use a standard MacBook Air (M2 chip, 16 GB RAM) for benchmarking. All runs used $n_{\mathrm{live}}=100$ and $n_{\mathrm{delete}}=50$.
Table 14: Wall-clock runtimes (seconds) for ARIMA grid searches of increasing size on a synthetic time series of $N=1000$ data points, evaluated on a standard consumer laptop (Apple M2 MacBook Air, 16 GB RAM). All runs used $n_{\mathrm{live}}=100$ and $n_{\mathrm{delete}}=50$. The timings are evaluated for different prior scales $\sigma^{\prime\prime}$ of the ARMA coefficients (Table 1).
Total Time (s)
Grid $\sigma^{\prime\prime}=1$ $\sigma^{\prime\prime}=0.8$ $\sigma^{\prime\prime}=0.6$
$2\times 2$ 71.12 72.57 70.79
$5\times 5$ 1046.20 1036.95 1027.45
$7\times 7$ 3473.27 3318.80 3276.66